of The location of a car y(t) is a function of time t as shown in the graph. Mark all statements that are true: the
media%2F002%2F002e6914-768f-48dc-ad49-dc
I’m either missing one or I checked the wrong box or maybe both.
Answer #1

dy/dt=v(t) dv/dt =a(t)

If you carefully observe the graph at t = 0, dy/dt=0 that means velocity is zero so the first option needs to be ticked car stopped and begins to move .At t=1 the y is a quadratic function that means dy/dt and d^2y/dt^2 exists .so it is accelerating at t=1 ,At t=2 the y(t) graph is a straight line that dy/dt is constant value and d^2y/dt^2 is zero ,so At t=2 car is moving at a steady speed .

Finally options to be ticked are 1 ,5, 9.